OpenStudy (anonymous):

How do you complete the square of f(x)=-1/2x^2+4x

OpenStudy (anonymous):

First factor out the \(-1/2\). This gets you: \[ -2 \cdot f(x)=x^2-8x \]Notice how: \[ (x+a)^2 = x^2+2ax+a^2 \]In this case \(2a=-8\) thus \(a=-4\). So we want: \[ x^2-8x+16 \]We do so by adding to both sides:\[ -2 \cdot + 16 f(x)=x^2-8x+16 = (x-4)^2 \]

OpenStudy (anonymous):

\[ -2 \cdot f(x)+16=x^2-8x+16 = (x-4)^2 \]Then we solve for \(f(x)\), keeping our square but undoing other the stuff we did.\[ \begin{array}{rcl} -2\cdot f(x) &=& (x-4)^2-16 \\ f(x) &=& -\frac{1}{2}(x-4)^2+8 \end{array} \]

OpenStudy (anonymous):

Thank you!!!